complete the equation in your answer booklet for the decay of phosphorus-30 by writing a notation for the missing product.


1
Expert's answer
2022-06-06T17:05:07-0400

Phosphorus-30 decays by positron emission.

We need to start by writing the parent element symbol (phosphorus is P). The number going on top of the symbol is the mass number(30) and the number on the bottom is the atomic number (15). Atomic number can be easily found on the periodic table at the bottom of the element. The mass number is given to us by the question. It is the number after the element name.

Since it is an emission, we must put the nuclear particle on the right side of the equation. Positron emission means we need to put positron symbol on the right side.

The equation is already set up for us in the answer booklet and all we need to do is to find the missing product.

How to balance nuclear equations? Sum of mass numbers on the left must equal to the sum of mass numbers on the right. Likewise, sum of atomic numbers on the left must equal to the sum go atomic numbers on the right.

Therefore, Mass numbers: 30=0+x, x=30 (mass number for the missing product)

Atomic numbers: 15=1+x, x=14 (atomic number for the missing element)

Now, all is left is to find the correct symbol for the element. For this purpose, we must use the atomic number (always!!!). Looking at the periodic table, we see that atomic number of 14 belongs to Si.

"\begin{smallmatrix}
 30 & \\
 15 & 
\end{smallmatrix}\large P⇒\begin{smallmatrix}
 0 & \\
 +1 & 
\end{smallmatrix} \large e 
+ \begin{smallmatrix}
 30 & \\
 14 & 
\end{smallmatrix}Si"